Understanding risk:
  • Part A: Stock A has a standard deviation of 10% and an expected return of 8%. Stock B has a standard deviation of 15% and an expected return of 11%. A client wants to know which stock has a better risk-return profile. How would you answer her?
  • Part B: Stock C has a standard deviation of 20% and a beta of 1.20. Stock D has a standard deviation of 16% and a beta of 1.44. A client wants to know which stock is riskier. How would you answer her?
Answer #1

1. a. Return of stock A = 8%
Return of stock B= 11%

b. Standard deviation of stock A = 10%
Standard deviation of stock B = 15%

D. Coefficient of variation (CV)= standard deviation/Return.

CV of A = 10/8= 1.25
CV of B = 15/11= 1.36

Coefficient of variation the lower the better. Stock A provides a better risk-return profile.

2. Stock C:
the standard deviation of 20%
the beta of 1.20.

Stock D:
the standard deviation of 16%  
the beta of 1.44.

Total risk= systematic risk+unsystematic risk
Stock D has more systematic risk as its beta is higher.
Stock C is more total risk as its standard deviation is higher.
Stock C is more risker due to its total risk.